Evaluate the "\oint"(12z-7)/(z-1)2(2z+3) dz where c is the circle |z+i|= Sqrt3

1
Expert's answer
2021-09-06T15:23:04-0400

Singular points of the integrand:

"{z_1} = 1" - second order pole,"\,\,{z_2} = - \frac{3}{2}" - simple pole.

Only a point "{z_1} = 1" falls inside the circle "|z + i| = \sqrt 3"



Let's find the residue at this point

"\mathop {res}\limits_{z = 1} f(z) = \mathop {\lim }\limits_{z \to 1} {\left( {\frac{{12z - 7}}{{{{(z - 1)}^2}\left( {2z + 3} \right)}}{{\left( {z - 1} \right)}^2}} \right)^\prime } = \mathop {\lim }\limits_{z \to 1} {\left( {\frac{{12z - 7}}{{2z + 3}}} \right)^\prime } = \mathop {\lim }\limits_{z \to 1} \frac{{12(2z + 3) - (12z - 7) \cdot 2}}{{{{\left( {2z + 3} \right)}^2}}} = \frac{{12 \cdot 5 - 2\left( {12 - 7} \right)}}{{{5^2}}} = 2"

Then

"\int\limits_{|z + i| = \sqrt 3 } {\frac{{12z - 7}}{{{{(z - 1)}^2}\left( {2z + 3} \right)}}dz} = 2\pi i\sum {\mathop {res}\limits_{z = {z_k}} } f(z) = 2\pi i \cdot 2 = 4\pi i"

Answer:"4\pi i"